OpenStudy (anonymous):

Find the slope of the line (-3,1) (2,-3) M=?

OpenStudy (anonymous):

Use the standard relationship\[m=\frac{y_2-y_1}{x_2-x_1}\]where your two given points are (x1,y1), and (x2,y2).

OpenStudy (anonymous):

OK, LIke this m=-3-1/2-(-3)=-4/5

OpenStudy (anonymous):

Looks good to me.

OpenStudy (anonymous):